Mark is drafting his fantasy football team. Right now, he can choose from 3 quarterbacks, 2 wide-receivers, and 5 linebackers. What is the probability that he will select a quarterback and then a wide-receiver?
1
Expert's answer
2018-05-11T08:34:08-0400
Mark is drafting his fantasy football team. Right now, he can choose from 3 quarterbacks, 2 wide-receivers, and 5 linebackers.
The probability that he will select a quarterback and then a wide-receiver is
P(A and B) = 3/10*2/9 = 1/18.